• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Snyder

    Log entry time 04:01:41 on July20,2004

    Entry number 129012

    Followups:

    keyword=Compton run problem

    During Compton run 8773, the cavity value shot up as indicated in the plot below. This run was stopped, coda was stopped, rebooted, then restarted. But, run 8775 seems to have the problem that the laser won't come on as shown in the second plot. This was also noticed in run 8773. Our best guess is simply that the laser isn't on.


    FIGURE 1

    FIGURE 2